3(2x – 3) = 4(2x + 4)
3(2x – 3) = 4(2x + 4)
3(2x – 3) = 4(2x + 4)
3(2x – 3) = 4(2x + 4)
3(2x – 3) = 4(2x + 4)
3(2x – 3) = 4(2x + 4)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3(2x – 3) = 4(2x + 4)

6x – 9 = 8x + 16

6x – 8x = 16 + 9

-2x = 25

x = -25/2

Thanks